OpenStudy (anonymous):

't' is proportional to the square root of 'd'. if t=12, d=4. a. Find a formula for 't' in terms of 'd' . b. Calculate the value of 't' when 'd' =9 ! :) if you answer it, u r my god....

OpenStudy (saifoo.khan):

\[\Huge t = k \sqrt{d}\]

OpenStudy (saifoo.khan):

\[12 = k \sqrt{4} \to 12 = k \times 2 \to k =\frac{12}{2} \to k = 6\]

OpenStudy (saifoo.khan):

\[t = 6\sqrt{d}\]

OpenStudy (saifoo.khan):

a) \[\huge d = \frac{t^2}{36}\]

OpenStudy (saifoo.khan):

i think u can solve b.

OpenStudy (anonymous):

thnks dude! :)

OpenStudy (saifoo.khan):

welks Suvi. =)